On January 1, 2018, Pomegranate Company acquired 90% of the voting stock of Starfruit Company for $91,700,000 in cash. The fair value of the noncontrolling interest in Starfruit at the date of acquisition was $6,300,000. Starfruit’s book value was $13,000,000 at the date of acquisition. Starfruit’s assets and liabilities were reported on its books at values approximating fair value, except its plant and equipment (10-year life, straight-line) was overvalued by $25,000,000. Starfruit Company had previously unreported intangible assets, with a market value of $40,000,000 and 5-year life, straight-line, which were capitalized following GAAP.

At the date of acquisition, consolidation eliminating entry (R) credits the noncontrolling interest in Starfruit Company in the amount of

$1,500,000

$5,000,000

$6,300,000

$8,500,000
Answer #1

Total fair value = 91,700,000 + 6,300,000 = 98,000,000

Fair value of identifiable assets = 13,000,000 + 40,000,000 - 25,000,000 = 28,000,000

Total goodwill = 98,000,000 - 28,000,000 = 70,000,000

Pomegranate goodwill = 91,700,000 - (28,000,000 * 90/100) = 66,500,000

Goodwill to NCI =70,000,000 - 66,500,000 = 3,500,000

NCI in Starfruit Company = [[40,000,000 - 25,000,000 ]*10/100] + 3,500,000 = 5,000,000